### v1

@@ 6 Conclusion
In this paper, based on a two-dimensional sentence representation, we propose a bi-consolidating model to address the semantic overlapping problem. Our model consists of a local consolidation component and a global consolidation component. The first component uses a pixel difference convolution to enhance semantic information of a possible triple representation of adjacent regions and mitigate noise in neighbouring neighbours. The second component strengthens the triple representation based a channel attention and a spatial attention, which has the advantage to learn remote semantic dependencies in a sentence. The bi-consolidating model simultaneously reinforce the local and global semantic features relevant to each relation triple. Our experiments demonstrate that on many benchmark datasets, the bi-consolidating model exceeds the performance of the previous state-of-the-art models. In-depth analyses are carried out to comprehend the influence of the components of the bi-consolidating model on performance and to validate their significance. Future research can focus on developing this model as a generalized framework to support various related multi-objective NLP tasks.
@@ Acknowledgements
This work is supported by National Natural Science Foundation of China under Grant No. 62066008 and No. 62066007, National Key R&D Program of China under Grant No. 2023YFC3304500 and the Major Science and Technology Projects of Guizhou Province under Grant [2024]003.

### v2

@@ 6 Conclusion
In this paper, based on a two-dimensional sentence representation, we propose a bi-consolidating model to address the semantic overlapping problem. Our model consists of a local consolidation component and a global consolidation component. The first component uses a pixel difference convolution to enhance semantic information of a possible triple representation of adjacent regions and mitigate noise in neighbouring neighbours. The second component strengthens the triple representation based a channel attention and a spatial attention, which has the advantage to learn remote semantic dependencies in a sentence. The bi-consolidating model simultaneously reinforce the local and global semantic features relevant to each relation triple. Our experiments demonstrate that on many benchmark datasets, the bi-consolidating model competitive performance. In-depth analyses are carried out to comprehend the influence of the components of the bi-consolidating model on performance and to validate their significance. Future research can focus on developing this model as a generalized framework to support various related multi-objective NLP tasks.
@@ Acknowledgements
This work is supported by National Key R&D Program of China under Grant No. 2023YFC3304500, the Major Science and Technology Projects of Guizhou Province under Grant [2024]003 and National Natural Science Foundation of China under Grant No. 62066008 and No. 62066007.